OpenStudy (anonymous):

The function f(x) = 15(2)x represents the growth of a frog population every year in a remote swamp. Elizabeth wants to manipulate the formula to an equivalent form that calculates every half-year, not every year. Which function is correct for Elizabeth's purposes?

OpenStudy (anonymous):

f(x) = 15(2 to the one half power)2x f(x) = 15(22)the x over 2 power f(x) = fifteen halves(2)x f(x) = 30(2)x these are the answer choices

OpenStudy (anonymous):

if x is years, the one half year is \(\frac{x}{2}\) right?

OpenStudy (anonymous):

so the second choice

OpenStudy (anonymous):

i can't read those

OpenStudy (jchick):

\(\large 15(2)^x = 15(2^{\frac{1}{2}})^{2x}\)